PLEASEE ANSWER ASAPPPP
Calculate the molar mass of of an unknown gas that weighs 4.47 grams in a 1.5L volume at 733 mmg and 46°C
Answer #1

Volume of the gas (V) = 1.5 L

Pressure (P) = 733 mmHg

Temperature (T) = 46 oC = (46 + 273) K = 319 K

Moles of the gas (n) = m/M

where,

m = mass of the gas = 4.47 g

M = molar mass of the gas

Universal gas constant (R) = 62.36 mmHg.L/mol.K

Now, substituting these values in the ideal gas equation to get,

PV = nRT

or, PV = mRT/M

or, 733 mmHg x 1.5 L = (4.47 g x 62.36 mmHg.L/mol.K x 319 K)/M

or, M = 80.9 g/mol

Hence, the molar mass of the gas = 80.9 g/mol